OpenStudy (anonymous):

together lauren and aaron sold 456 raffle tickets. lauren sold 24 more than aron. how many did lauren sell

OpenStudy (campbell_st):

let Aarons ticket sales = x and Lauren's will be x + 24 so the total sales x + x + 24 = 456 solve for x

OpenStudy (anonymous):

campbell whats the nxt step

OpenStudy (anonymous):

This is a systems of equations question. L is Lauren, A is Aron. L + A = 456 L = A + 24 Rearrange so you can solve for L. A = 456 - L L = A + 24 Substitute to solve for L. L = (456 - L) + 24 Simplify. L = 480 - L Simplify. 2L = 480 Simplify. L = 240. Lauren sold 240 tickets.

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

well its 2x + 24 = 456